A company has net sales of $750,000, cost of goods sold is $500,000, net income is $20,000. What are the company’s gross margin expense and operating expense?
Answer #1

Answers:

Gross MArgin = $250,000

Operating Expenses = $ 230,000

Gross MArgin = Net sales - Cost of Goods Sold = $750,000-500,000 = $ 250,000

Operating Expenses = Gross MArgin - Net income = 250,000-20,000 = $230,000